OpenStudy (anonymous):

Integrate (x^2+25)^-1 from 5 to infinity. How would I do this WITHOUT using tables?

OpenStudy (anonymous):

I know that this is equivalent to inverse tangent, but is that using a table?

OpenStudy (experimentx):

then, let x = 5 tan\theta

OpenStudy (experimentx):

use trigonometric substitution, you will arrive at the same conclusion.

OpenStudy (anonymous):

you can always substitute... c in the denominator that x2 +25 can be reduced to y2+1just by putting x= 5y.... but also c that u can put x=5 * tan(theta) which would reduce the lower term to just 1 term i.e sec2(theta) and u can integrate very easily...

OpenStudy (anonymous):

Using inverse tangent (aka arctan) is not considered using tables.

OpenStudy (anonymous):

\[\frac{\pi}{20}\]